# Adaptive-TMLE for the Average Treatment Effect based on Randomized Controlled Trial Augmented with Real-World Data

### v1

@@ 4.4 TMLE of the projection parameter $\Psi^{\#}_{{\cal M}_{w,2,n}}$
while for inference for the original target parameter $\Psi^{\#}(P_{0})$, we also use that $\Psi^{\#}_{{\cal M}_{w,2,n}}(P_{0})-\Psi^{\#}(P_{0})=o_{P}(n^{-1/2})$, under reasonable regularity conditions as discussed in [adml].
In this section, we discuss the estimation of the pooled-ATE estimand $\tilde{\Psi}(P_{0})$, which follows a very similar procedure as in the bias estimand $\Psi^{\#}(P_{0})$ case we described above. Except now, we will construct a data-adaptive working model for the conditional effect of the treatment $A$ on the outcome $Y$ instead of $S$ on $Y$.
5.1 Semiparametric regression working model for the conditional effect of the treatment
Let ${\cal Q}_{w}^{r}=\{\theta(W)+AT_{\beta}(W):\theta,\beta\}$, where $T_{\beta}(W)=\sum_{j}\beta(j)\phi_{j}(W)$. Let $\bar{Q}_{w,P}^{r}=\arg\min_{\bar{Q}^{r}\in{\cal Q}_{w}^{r}}P(\bar{Q}_{P}^{r}-\bar{Q}^{r})^{2}$, where $\bar{Q}_{P}^{r}=E_{P}(Y\mid W,A)$.
Lemma 6.
Recall $T_{P}(W)=E_{P}(Y\mid W,A=1)-E_{P}(Y\mid W,A=0)$. We have that $\bar{Q}_{w,P}^{r}=E_{P}(Y\mid W)+\sum_{j}(A-g_{P}(1\mid W))\beta_{P}(j)\phi_{j}(W)$, where
$\beta(P)=\arg\min_{\beta}E_{P}g_{P}(1-g_{P})(1\mid W)\left(T_{P}(W)-\sum_{j}\beta_{j}\phi_{j}(W)\right)^{2}.$
Thus, $\sum_{j}\beta_{P}(j)\phi_{j}$ represents the projection of $T_{P}(W)$ onto $\{\sum_{j}\beta_{j}\phi_{j}:\beta\}$ with respect to a weighted $L^{2}$-norm with weights $g_{P}(1-g_{P})(1\mid W)$. We also have
$\beta(P)=\arg\min_{\beta}E_{P}\left(\bar{Q}_{P}^{r}-(A-g_{P}(1\mid W))\sum_{j}\beta_{j}\phi_{j}(W)\right)^{2},$
and thereby
$\beta(P)=\arg\min_{\beta}E_{P}\left(Y-(A-g_{P}(1\mid W))\sum_{j}\beta_{j}\phi_{j}(W)\right)^{2}.$
Here one could replace $Y$ by $Y-E_{P}(Y\mid W)$ as well.
This also shows that
$T_{P}=\arg\min_{T}E_{P}(Y-E_{P}(Y\mid W)-(A-g_{P}(1\mid W))T(W))^{2}.$
Therefore, we can use the following loss function for learning $T_{P}$:
$L_{\tilde{\theta}^{r}_{P},g_{P}}(T)=(Y-\tilde{\theta}^{r}_{P}(W)-(A-g_{P}(1\mid W))T(W))^{2},$
where this loss function is indexed by the nuisance parameters $\tilde{\theta}^{r}_{P}(W)=E_{P}(Y\mid W)$ and $g_{P}(1\mid W)=P(A=1\mid W)$.
The proof can be found in Appendix A.
5.2 Canonical gradient of $\tilde{\Psi}_{\mathcal{M}_{w,1}}$ at $P$
First, we find the canonical gradient of the $\beta(P)$-component.
Lemma 7.
Let $\beta(P)$ be defined as in Lemma 6 on a nonparametric model. The canonical gradient of $\beta$ at $P$ is given by
$D_{\beta,P}^{r}=I_{P}^{-1}(A-g_{P}(1\mid W)){\bf\phi}(W)\left(Y-\tilde{\theta}_{P}^{r}(W)-(A-g_{P}(1\mid W))\sum_{j}\beta_{P}(j)\phi_{j}(W)\right),$
where $\tilde{\theta}_{P}^{r}=E_{P}(Y\mid W)$, $g_{P}(1\mid W)=P(A=1\mid W)$ and $I_{P}=E_{P}g_{P}(1-g_{P})(1\mid W){\bf\phi}{\bf\phi}^{\top}(W)$. By Lemma 6, this can also be written as
$D_{\beta,P}^{r}=I_{P}^{-1}(A-g_{P}(1\mid W)){\bf\phi}(W)(Y-\bar{Q}_{w,P}^{r}(W,A)).$
The proof can be found in Appendix A. The next lemma presents the canonical gradient of the projection parameter $\tilde{\Psi}_{\mathcal{M}_{w,1}}$ at $P$.
Lemma 8.
Let $\tilde{\Psi}_{{\cal M}_{w,1}}(P)=E_{P}T_{w,\beta(P)}=E_{P}(\bar{Q}_{w,P}^{r}(W,1)-\bar{Q}_{w,P}^{r}(W,0))$, while $\tilde{\Psi}(P)=E_{P}T_{P}=E_{P}(\bar{Q}_{P}^{r}(W,1)-\bar{Q}_{P}^{r}(W,0))$, where $\bar{Q}_{w,P}^{r}=\tilde{\theta}_{P}^{r}(W)+AT_{\beta(P)}$, $\beta(P)=\arg\min_{\beta}E_{P}g_{P}(1-g_{P})(1\mid W)(T_{P}(W)-T_{\beta}(W))$, and $T_{\beta}(W)=\sum_{j}\beta(j)\phi_{j}(W)$. We have that the canonical gradient of $\tilde{\Psi}_{{\cal M}_{w,1}}$ at $P$ is given by:
$D_{{\cal M}_{w,1},\tilde{\Psi},P}=T_{w,\beta(P)}-E_{P}T_{w,\beta(P)}+\sum_{j}D^{r}_{\beta,P,j}(O)E_{P}\phi_{j}(W),$
where
$D_{\beta,P}^{r}=(A-g_{P}(1\mid W))I_{P}^{-1}{\bf\phi}(W)(Y-\tilde{\theta}_{P}^{r}(W)-(A-g_{P}(1\mid W))\sum_{j}\beta_{P}(j)\phi_{j}(W)),$
with $D_{\beta,P,j}^{r}$ denoting the $j$-th component of $D_{\beta,P}^{r}$.
The proof can be found in Appendix A.
5.3 TMLE of the projection parameter $\tilde{\Psi}_{{\cal M}_{w,1,n}}$
To construct a TMLE for the pooled-ATE projection parameter, we first obtain initial estimators $g_{n}$ and $\tilde{\theta}_{n}^{r}$ of $g_{0}$ and $\tilde{\theta}_{0}^{r}$, respectively. If we use a relaxed-HAL for learning the working model for the conditional effect of $A$ on $Y$, then we have $\beta_{n}=\arg\min_{\beta}P_{n}L_{\tilde{\theta}_{n}^{r},g_{n}}(T_{w,n,\beta})$, which solves $P_{n}D_{\mathcal{M}_{w,1,n},\beta,\tilde{\theta}_{n}^{r},g_{n},\beta_{n}}=0$. We estimate $P_{W,0}$ with the empirical measure of $W$. We have then solved $P_{n}D_{{\cal M}_{w,1,n},\tilde{\Psi},\beta_{n},\tilde{\theta}_{n}^{r},g_{n}}=0$. The TMLE is now the plug-in estimator $\tilde{\Psi}_{\mathcal{M}_{w,1,n}}(P_{n}^{*})\equiv\tilde{\Psi}_{{\cal M}_{w,1,n}}(P_{W,n},\beta_{n})$.
@@ 5.4 Corresponding projection parameter of the original parameter $\Psi(P_{0})$ and its A-TMLE
Let $\mathcal{M}_{w,n}=\mathcal{M}_{w,1,n}\times\mathcal{M}_{w,2,n}$, we can now also define the projection parameter of our target parameter $\Psi$ as
$\Psi_{\mathcal{M}_{w,n}}(P)=\tilde{\Psi}_{\mathcal{M}_{w,1,n}}(P)-\Psi^{\#}_{{\cal M}_{w,2,n}}(P)=\tilde{\Psi}(\Pi_{{\cal M}_{w,1,n}}(P))-\Psi^{\#}(\Pi_{{\cal M}_{w,2,n}}(P)).$
@@
Note that we are using different working models for $\Psi^{\#}$ and $\tilde{\Psi}$, since $\tilde{\Psi}_{\mathcal{M}_{w,1,n}}$ depends on $P$ through $P_{W},T_{w,n,P}$ and $\Pi_{P}$, while $\Psi^{\#}_{{\cal M}_{w,2,n}}$ depends on $P$ through $P_{W},\tau_{w,n,P}$ and $\Pi_{P}$. The final A-TMLE estimator is obtained by plugging in the TMLEs for the projection parameters of $\tilde{\Psi}$ and $\Psi^{\#}$:
$\Psi(P_{n}^{*})=\Psi_{\mathcal{M}_{w,n}}(P_{n}^{*})=\tilde{\Psi}_{\mathcal{M}_{w,1,n}}(P_{n}^{*})-\Psi^{\#}_{\mathcal{M}_{w,2,n}}(P_{n}^{*}).$
6 Implementation of A-TMLE
In this section, we describe the algorithmic implementation of A-TMLE and discuss strategies for data-adaptively learning the working models. For illustration purposes, we will add some context by considering a simple synthetic study to which A-TMLE could be applied. Consider a study where researchers aim to assess the impact of a new medication on systolic blood pressure (SBP), a continuous outcome variable, versus that of standard-of-care (control arm). Suppose that the true ATE of the medication is 10 mmHg, that is, had patients taken the medication, their SBP would drop by 10 mmHg. Further suppose we adopt a hybrid study design combining an RCT with RWD from electronic health records. Say researchers identified a set of confounders including age, body mass index (BMI), smoking status, health literacy, and socio-economic status. However, health literacy and socio-economic status are not captured in neither RCT nor the RWD. Although the RCT is free from confounding bias due to the treatment randomization, in the RWD, the unmeasured confounders affect both the treatment and the SBP. As a result, a direct pooling of the two data sources would produce a biased estimate of the true ATE. The data-generating process can be found in Appendix B. Investigators could apply A-TMLE in this setting to improve the accuracy and precision of the ATE estimate. At a high-level, A-TMLE constructs estimators for the pooled-ATE and bias estimand separately. In this context, to estimate the pooled-ATE, data from both the RCT and RWD are combined, as described in Section 5. Due to the unmeasured confounders, this direct pooling is likely to be biased. To correct for this, the next step involves estimating the bias following the procedures outlined in Section 4. Intuitively, the bias can be decomposed into its impact on the treatment and control arms. As an example, for the treatment arm, we consider how trial enrollment under the same treatment regime might alter a subject’s SBP. This subproblem resembles an ATE estimation problem, with trial participation becoming the treatment variable. Here, some form of weighting by the probability of trial participation happens within the bias term, as detailed in Section 3. The final ATE estimate is obtained by subtracting the bias estimate from the pooled-ATE. The complete A-TMLE steps are summarized in Algorithm 1.
Algorithm 1 Adaptive-TMLE
1:Initial estimator $\Psi_{\mathcal{M}_{w,2,n}}^{\#}(P_{n})$ of $\Psi_{\mathcal{M}_{w,2,n}}^{\#}(P_{0})$:
2:     Obtain nuisance estimate $\Pi_{n}$ of $\Pi_{0}(1\mid W,A)\equiv P(S=1\mid W,A)$;
3:     Obtain nuisance estimate $\theta_{n}$ of $\theta_{0}(W,A)\equiv E(Y\mid W,A)$;
4:     Compute pseudo outcome $Y_{\Psi^{\#},\text{pseudo}}\equiv(Y-\theta_{n})/(S-\Pi_{n})$ and weight $\mathcal{W}_{\Psi^{\#},\text{pseudo}}\equiv(S-\Pi_{n})^{2}$;
5:     Fit relaxed-HAL using $(W,A)$ as covariates, $Y_{\Psi^{\#},\text{pseudo}}$ as outcome, and $\mathcal{W}_{\Psi^{\#},\text{pseudo}}$ as weight;
6:     Get initial estimate $\Psi_{\mathcal{M}_{w,2,n}}^{\#}(P_{n})\equiv 1/n\sum_{i=1}^{n}(\Pi_{n}(0\mid W_{i},0)\tau_{n}(W_{i},0)-\Pi_{n}(0\mid W_{i},1)\tau_{n}(W_{i},1))$;
7:Initial estimator $\tilde{\Psi}_{\mathcal{M}_{w,1,n}}(P_{n})$ of $\tilde{\Psi}_{\mathcal{M}_{w,1,n}}(P_{0})$:
8:     Obtain nuisance estimate $g_{n}$ of $g_{0}(1\mid W)\equiv P(A=1\mid W)$;
9:     Obtain nuisance estimate $\tilde{\theta}_{n}$ of $\tilde{\theta}_{0}(W)\equiv E(Y\mid W)$;
10:     Define pseudo outcome $Y_{\tilde{\Psi},\text{pseudo}}\equiv(Y-\theta_{n})/(A-g_{n})$ and weight $\mathcal{W}_{\tilde{\Psi},\text{pseudo}}\equiv(A-g_{n})^{2}$;
11:     Fit relaxed-HAL using $W$ as covariates, $Y_{\tilde{\Psi},\text{pseudo}}$ as outcome, and $\mathcal{W}_{\tilde{\Psi},\text{pseudo}}$ as weight;
12:     Get initial estimate $\tilde{\Psi}_{\mathcal{M}_{w,1,n}}(P_{n})\equiv 1/n\sum_{i=1}^{n}(T_{n}(W_{i}))\equiv\tilde{\Psi}_{\mathcal{M}_{w,1,n}}(P_{n}^{*})$;
13:Additional TMLE targeting for $\Pi$-component:
14:     Compute clever covariate $C(g_{n},\beta_{n})=I(A=1)/g_{n}(1\mid W)\tau_{n}(W,1)-I(A=0)/g_{n}(0\mid W)\tau_{n}(W,0)$;
15:     Use $C(g_{n},\beta_{n})$ to perform a TMLE targeting and obtain an updated estimate $\Pi^{*}_{n}$ of $\Pi_{0}$;
16:     Get updated estimate $\Psi_{\mathcal{M}_{w,2,n}}^{\#}(P_{n}^{*})\equiv 1/n\sum_{i=1}^{n}(\Pi_{n}^{*}(0\mid W_{i},0)\tau_{n}(W_{i},0)-\Pi_{n}^{*}(0\mid W_{i},1)\tau_{n}(W_{i},1))$;
17:Obtain the final estimate $\Psi_{\mathcal{M}_{w,n}}(P_{n}^{*})=\tilde{\Psi}_{\mathcal{M}_{w,1,n}}(P_{n}^{*})-\Psi^{\#}_{{\cal M}_{w,2,n}}(P_{n}^{*})$;
18:Compute the 95% confidence interval given by $\Psi_{\mathcal{M}_{w,n}}(P_{n}^{*})\pm 1.96\sqrt{P_{n}(D_{\mathcal{M}_{w,1,n},\tilde{\Psi}}-D_{\mathcal{M}_{w,2,n},\Psi^{\#}})^{2}/n}$.
Figure 1: Results for the synthetic data. Point estimates and 95% confidence intervals for the reduction in systolic blood pressure comparing using RCT data only and A-TMLE.
We make the following remarks regarding the algorithm. To estimate nuisance parameters, one might use either highly adaptive lasso (HAL) [benkeser_hal_2016, hejazi_hal_2020], or super learner [vdl_sl_2007, vdl_handbook_2020]. For super learner in particular, one could incorporate a rich library of flexible machine learning algorithms including HAL. Specifically for learning the two working models, we recommend using relaxed-HAL. This approach involves initially fitting a HAL, followed by applying an ordinary least squares regression to the spline basis functions that exhibit non-zero coefficients. Because relaxed-HAL is an MLE, the empirical mean of the $\beta$-component of the canonical gradient is automatically solved. Notably, HAL converges in loss-based dissimilarity at a rate of $n^{-1/3}(\log n)^{2(d-1)/3}$ [bibaut_rate_2019]. Importantly, this rate depends on the dimensionality $d$ only via a power of the $\log n$-factor thus is essentially dimension-free. This allows the learned working model to approximate the truth well. In Appendix D, we also consider several alternative approaches for constructing data-adaptive working models, including learning the working model on independent data sets, and using deep learning architectures to generate dimension reductions of data sets. To complement the methods discussed in Appendix D, we also propose and analyze a cross-validated version of A-TMLE in Appendix C to address concerns regarding potential overfitting of the working models as a result of being too data-adaptive. In addition, as suggested in the Introduction, opting for a slightly larger model than indicated by cross-validation could reduce bias. Techniques like undersmoothing as described in [undersmooth] are worth considering. Furthermore, to ensure robustness in model selection, one might enforce a minimal working model based on domain knowledge, specifying essential covariates that must be included. For instance, using HAL to develop the working model could mean exempting certain critical covariates from penalization to guarantee their presence in the final model. This option is available in standard lasso software like the ‘glmnet’ R package [friedman_lasso_2010].
Algorithm 1 has been implemented in the R package ‘atmle’ (available at https://github.com/tq21/atmle). Going back to the synthetic SBP example, we run the implemented algorithm on the simulated data. As we see in Figure 1, the confidence interval produced by A-TMLE is significantly narrower than that obtained from RCT data alone. The red dashed line indicates the true ATE.
@@ 7 Asymptotic Super-Efficiency of A-TMLE
In the previous section, we provided detailed steps and strategies for implementing an A-TMLE for our target estimand. In this section, we examine the theoretical properties of A-TMLE. Let $\mathcal{M}_{w,n}=\mathcal{M}_{w,1,n}\times\mathcal{M}_{w,2,n}$. We establish the asymptotic super-efficiency of the A-TMLE $\Psi_{\mathcal{M}_{w,n}}(P_{n}^{*})$ as an estimator of $\Psi(P_{0})$, under specified conditions. This is completely analogue to [adml], but is presented to make this article self-contained. It follows the proof of asymptotic efficiency for TMLE applied to $\Psi_{\mathcal{M}_{w,n}}(P_{n}^{*})-\Psi_{\mathcal{M}_{w,n}}(P_{0})$, but with the additional work 1) to deal with the data dependent efficient influence curve in its resulting expansion and 2) to establish that the bias $\Psi_{\mathcal{M}_{w,n}}(P_{0})-\Psi(P_{0})$ is second order. The conditions will be similar to the empirical process and second-order remainder conditions needed for analyzing a TMLE with an additional condition on the data-adaptive working model $\mathcal{M}_{w,n}$ with respect to approximating $P_{0}$.
Let $D_{\mathcal{M}_{w,n},P}$ be the canonical gradient of $\Psi_{\mathcal{M}_{w,n}}$ at $P$. Let $R_{\mathcal{M}_{w,n}}(P,P_{0})\equiv\Psi_{\mathcal{M}_{w,n}}(P)-\Psi_{\mathcal{M}_{w,n}}(P_{0})+P_{0}D_{\mathcal{M}_{w,n},P}$ be the exact remainder. Let $P_{n}^{*}\in{\cal M}_{w,n}$ be an estimator of $P_{0}$ with $P_{n}D_{{\cal M}_{w,n},P_{n}^{*}}=o_{P}(n^{-1/2})$. An MLE over ${\cal M}_{w,n}$ or a TMLE starting with an initial estimator $P_{n}^{0}\in{\cal M}_{w,n}$ targeting $\Psi_{{\cal M}_{w,n}}(P_{0})$ satisfies this efficient score equation condition. Then, by definition of the exact remainder, we have
$\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi_{{\cal M}_{w,n}}(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w,n},P_{n}^{*}}+R_{{\cal M}_{w,n}}(P_{n}^{*},P_{0})+o_{P}(n^{-1/2}).$
We assume that ${\cal M}_{w,n}$ approximates an oracle model ${\cal M}_{0}$ that contains $P_{0}$ so that
$\Psi_{{\cal M}_{w,n}}(P_{0})-\Psi(P_{0})=o_{P}(n^{-1/2}).$
Let $\Psi_{{\cal M}_{0}}(P)=\Psi(\Pi(P\mid{\cal M}_{0}))$, i.e. the projection of $P$ onto $\mathcal{M}_{0}$ applied to $\Psi$. As shown in [adml], this does not represent an unreasonable condition. For example, if $P_{0,n}\equiv\Pi(P_{0}\mid{\cal M}_{w,n})\in{\cal M}_{0}$ with probability tending to 1, which holds if ${\cal M}_{w,n}\subset{\cal M}_{0}$, noting that $P_{0}\tilde{D}_{{\cal M}_{w,n},P_{0,n}}=0$ for any $\tilde{D}_{{\cal M}_{w,n},P_{0,n}}$ in the tangent space $T_{{\cal M}_{w,n}}(P_{0,n})$ of the working model at $P_{0,n}$ (due to $P_{0,n}$ be an MLE of $p\rightarrow P_{0}\log p$ over $P\in{\cal M}_{w,n}$). Then, we have
$\Psi_{{\cal M}_{w,n}}(P_{0})-\Psi(P_{0})$ $=\Psi(P_{0,n})-\Psi(P_{0})$
@@
$=(P_{0,n}-P_{0})\left\{D_{{\cal M}_{0},P_{0,n}}-\tilde{D}_{{\cal M}_{w,n},P_{0,n}}\right\}+R_{{\cal M}_{0}}(P_{0,n},P_{0}).$
@@
The exact remainder $R_{{\cal M}_{0}}(P_{0,n},P_{0})$ is second-order in $p_{0,n}-p_{0}$, so that, if ${\cal M}_{w,n}$ approximates $P_{0}$ at rate $n^{-1/4}$, then this will be $o_{P}(n^{-1/2})$. Regarding the leading term, we can select $\tilde{D}_{{\cal M}_{w,n},P_{0,n}}$ equal to the projection of $D_{{\cal M}_{0},P_{0,n}}$ onto $T_{{\cal M}_{w,n}}(P_{0,n})$ in $L^{2}(P_{0,n})$. Thus, under this condition $\Pi_{{\cal M}_{w,n}}(P_{0})\in{\cal M}_{0}$, using Cauchy-Schwarz inequality, we have that $\Psi_{{\cal M}_{w,n}}(P_{0})-\Psi(P_{0})$ is a second-order difference in two oracle approximation errors $\parallel p_{0,n}-p_{0}\parallel_{\mu}\parallel D_{{\cal M}_{0},P_{0,n}}-\Pi(D_{{\cal M}_{0},P_{0,n}}\mid T_{{\cal M}_{w,n}}(P_{0,n}))\parallel_{\mu}$, where $\mu$ is a dominating measure of $P_{0,n}$ and $P_{0}$. Moreover, one can weaken this condition by defining $\tilde{P}_{0,n}=\Pi(P_{0,n}\mid{\cal M}_{0})$, assuming the above second-order remainder with $P_{0,n}$ replaced by $\tilde{P}_{0,n}$ is $o_{P}(n^{-1/2})$, and assuming that $\Psi(\tilde{P}_{0,n})-\Psi(P_{0,n})=o_{P}(n^{-1/2})$, thereby only assuming that $P_{0,n}$ is close enough to ${\cal M}_{0}$ (instead of being an element of ${\cal M}_{0}$) defined by a distance induced by $\Psi$. So under this reasonable approximation condition on ${\cal M}_{w,n}$ with respect to an oracle model ${\cal M}_{0}$, we have
$\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w,n},P_{n}^{*}}+R_{{\cal M}_{w,n}}(P_{n}^{*},P_{0})+o_{P}(n^{-1/2}).$
@@
Under the condition that $P_{n}^{*}$ converges to $P_{0}$ at a fast enough rate (i..e, $o_{P}(n^{-1/4})$ so that $R_{{\cal M}_{w,n}}(P_{n}^{*},P_{0})=o_{P}(n^{-1/2})$, this yields
$\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w,n},P_{n}^{*}}+o_{P}(n^{-1/2}).$
@@
This condition would hold if we use HAL to construct the initial estimator of $P_{0}$ in the TMLE. Under the Donsker class condition that $D_{{\cal M}_{w,n},P_{n}^{*}}$ falls in a $P_{0}$-Donsker class with probability tending to 1, this implies already $\Psi_{{\cal M}_{w,n}}(P_{n}^{*})-\Psi(P_{0})=O_{P}(n^{-1/2})$, consistency at the parametric rate $n^{-1/2}$. Inspection of the canonical gradients for our projection parameters $\Psi_{{\cal M}_{w,n}}$ shows that this is not a stronger condition than the Donsker class condition $D_{\Psi,P_{n}^{*}}$ one would need for the TMLE of $\Psi(P_{0})$. The condition on the working model ${\cal M}_{w,n}$ would hold if the model ${\cal Q}_{w,n}$ falls with probability tending to one in a nice class such as the class of cádlág functions with a universal bound on the sectional variation norm.

### v2

@@ 5.3 TMLE of the projection parameter $\Psi^{\#}_{{\cal M}_{S,w}}$
while for inference for the original target parameter $\Psi^{\#}(P_{0})$, we use that $\Psi^{\#}_{{\cal M}_{S,w}}(P_{0})-\Psi^{\#}(P_{0})=o_{P}(n^{-1/2})$, under reasonable regularity conditions as discussed in [adml] and also the section that follows.
@@ 5.4 Corresponding projection parameter of the original parameter $\Psi(P_{0})$ and its A-TMLE
Let $\mathcal{M}_{w}=\mathcal{M}_{A,w}\times\mathcal{M}_{S,w}$, we can now also define the projection parameter of our target parameter $\Psi$ as
$\Psi_{\mathcal{M}_{w}}(P)=\tilde{\Psi}(\Pi_{{\cal M}_{A,w}}(P))-\Psi^{\#}(\Pi_{{\cal M}_{S,w}}(P))=\tilde{\Psi}_{\mathcal{M}_{A,w}}(P)-\Psi^{\#}_{{\cal M}_{S,w}}(P).$
@@
Note that we are using different working models for $\Psi^{\#}$ and $\tilde{\Psi}$, since $\tilde{\Psi}_{\mathcal{M}_{A,w}}$ depends on $P$ through $P_{W}$ and $\tau_{A,\beta_{P}}$, while $\Psi^{\#}_{{\cal M}_{S,w}}$ depends on $P$ through $P_{W},\tau_{S,\beta_{P}}$ and $\Pi_{P}$. The final A-TMLE estimator is obtained by plugging in the TMLEs for the projection parameters of $\tilde{\Psi}$ and $\Psi^{\#}$:
$\Psi(P_{n}^{\star})=\Psi_{\mathcal{M}_{w}}(P_{n}^{\star})=\tilde{\Psi}_{\mathcal{M}_{A,w}}(P_{n}^{\star})-\Psi^{\#}_{\mathcal{M}_{S,w}}(P_{n}^{\star}).$
We also have the option to estimate $\tilde{\Psi}(P_{0})$ with the regular TMLE and only use an A-TMLE for $\Psi^{\#}_{\mathcal{M}_{S,w}}(P_{0})$.
@@ 6 Asymptotic Super-Efficiency of A-TMLE
We now examine the theoretical properties of A-TMLE. Specifically, we establish the asymptotic super-efficiency of the A-TMLE $\Psi_{\mathcal{M}_{w}}(P_{n}^{\star})$ as an estimator of $\Psi(P_{0})$, under certain conditions. While this analysis closely mirrors the work in [adml], it is included here for completeness and to make this article self-contained. It follows the proof of asymptotic efficiency for TMLE applied to $\Psi_{\mathcal{M}_{w}}(P_{n}^{\star})-\Psi_{\mathcal{M}_{w}}(P_{0})$, but with additional work to 1) deal with the data-dependent efficient influence curve in its resulting expansion and 2) establish that the bias $\Psi_{\mathcal{M}_{w}}(P_{0})-\Psi(P_{0})$ is second-order. The required conditions are analogous to the empirical process and second-order remainder conditions necessary for establishing asymptotic efficiency of a standard TMLE, with an added condition on the data-adaptive working model $\mathcal{M}_{w}$ with respect to approximating the true data-generating distribution $P_{0}$.
Let $D_{\mathcal{M}_{w},P}$ be the canonical gradient of $\Psi_{\mathcal{M}_{w}}$ at $P$. Let $R_{\mathcal{M}_{w}}(P,P_{0})\equiv\Psi_{\mathcal{M}_{w}}(P)-\Psi_{\mathcal{M}_{w}}(P_{0})+P_{0}D_{\mathcal{M}_{w},P}$ be the exact remainder. Let $P_{n}^{\star}\in{\cal M}_{w}$ be an estimator of $P_{0}$ with $P_{n}D_{{\cal M}_{w},P_{n}^{\star}}=o_{P}(n^{-1/2})$. An MLE over ${\cal M}_{w}$ or a TMLE starting with an initial estimator $P_{n}^{0}\in{\cal M}_{w}$ targeting $\Psi_{{\cal M}_{w}}(P_{0})$ satisfies this efficient score equation condition. Then, by definition of the exact remainder, we have
$\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi_{{\cal M}_{w}}(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w},P_{n}^{\star}}+R_{{\cal M}_{w}}(P_{n}^{\star},P_{0})+o_{P}(n^{-1/2}).$
We assume that ${\cal M}_{w}$ approximates an oracle model ${\cal M}_{0}$ that contains $P_{0}$ so that
$\Psi_{{\cal M}_{w}}(P_{0})-\Psi_{\mathcal{M}_{0}}(P_{0})=o_{P}(n^{-1/2}),$
where $\Psi_{{\cal M}_{0}}(P)=\Psi(\Pi_{{\cal M}_{0}}(P))$, i.e., the projection of $P$ onto $\mathcal{M}_{0}$ applied to $\Psi$. As shown in [adml], this does not represent an unreasonable condition. For example, if $P_{0,n}\equiv\Pi_{\mathcal{M}_{w}}(P_{0})\in{\cal M}_{0}$ with probability tending to 1, which holds if ${\cal M}_{w}\subset{\cal M}_{0}$, noting that $P_{0}\tilde{D}_{{\cal M}_{w},P_{0,n}}=0$ for any $\tilde{D}_{{\cal M}_{w},P_{0,n}}$ in the tangent space $T_{{\cal M}_{w}}(P_{0,n})$ of the working model at $P_{0,n}$ (due to $P_{0,n}$ being an MLE of $p\rightarrow P_{0}\log p$ over $P\in{\cal M}_{w}$). Then, we have
$\Psi_{{\cal M}_{w}}(P_{0})-\Psi(P_{0})$ $=\Psi(P_{0,n})-\Psi(P_{0})$
@@
$=(P_{0,n}-P_{0})\left\{D_{{\cal M}_{0},P_{0,n}}-\tilde{D}_{{\cal M}_{w},P_{0,n}}\right\}+R_{{\cal M}_{0}}(P_{0,n},P_{0}).$
@@
The exact remainder $R_{{\cal M}_{0}}(P_{0,n},P_{0})$ is second-order in $p_{0,n}-p_{0}$, so that, if ${\cal M}_{w}$ approximates $P_{0}$ at rate $n^{-1/4}$, then this will be $o_{P}(n^{-1/2})$. Regarding the leading term, we can select $\tilde{D}_{{\cal M}_{w},P_{0,n}}$ equal to the projection of $D_{{\cal M}_{0},P_{0,n}}$ onto $T_{{\cal M}_{w}}(P_{0,n})$ in $L^{2}(P_{0,n})$. Thus, under the condition $\Pi_{{\cal M}_{w}}(P_{0})\in{\cal M}_{0}$, using Cauchy-Schwarz inequality, we have that $\Psi_{{\cal M}_{w}}(P_{0})-\Psi(P_{0})$ is a second-order difference in two oracle approximation errors $\parallel p_{0,n}-p_{0}\parallel_{\mu}\parallel D_{{\cal M}_{0},P_{0,n}}-\Pi_{T_{{\cal M}_{w}}(P_{0,n})}(D_{{\cal M}_{0},P_{0,n}})\parallel_{\mu}$, where $\mu$ is a dominating measure of $P_{0,n}$ and $P_{0}$. Moreover, one can weaken this condition by defining $\tilde{P}_{0,n}=\Pi_{{\cal M}_{0}}(P_{0,n})$, assuming the above second-order remainder with $P_{0,n}$ replaced by $\tilde{P}_{0,n}$ is $o_{P}(n^{-1/2})$, and assuming that $\Psi(\tilde{P}_{0,n})-\Psi(P_{0,n})=o_{P}(n^{-1/2})$, thereby only assuming that $P_{0,n}$ is close enough to ${\cal M}_{0}$ (instead of being an element of ${\cal M}_{0}$) defined by a distance induced by $\Psi$. So under this reasonable approximation condition on ${\cal M}_{w}$ with respect to an oracle model ${\cal M}_{0}$, we have
$\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w},P_{n}^{\star}}+R_{{\cal M}_{w}}(P_{n}^{\star},P_{0})+o_{P}(n^{-1/2}).$
@@
Under the condition that $P_{n}^{\star}$ converges to $P_{0}$ at a fast enough rate (i.e., $o_{P}(n^{-1/4})$ so that $R_{{\cal M}_{w}}(P_{n}^{\star},P_{0})=o_{P}(n^{-1/2})$, this yields
$\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w},P_{n}^{\star}}+o_{P}(n^{-1/2}).$
@@
This condition would hold if we use HAL to construct the initial estimator of $P_{0}$ in the TMLE. Under the Donsker class condition that $D_{{\cal M}_{w},P_{n}^{\star}}$ falls in a $P_{0}$-Donsker class with probability tending to 1, this implies already $\Psi_{{\cal M}_{w}}(P_{n}^{\star})-\Psi(P_{0})=o_{P}(n^{-1/2})$, consistency at the parametric rate $n^{-1/2}$. Inspection of the canonical gradient for our projection parameter $\Psi_{{\cal M}_{w}}$ shows that this is not a stronger condition than the Donsker class condition $D_{\Psi,P_{n}^{\star}}$ one would need for the TMLE of $\Psi(P_{0})$. The condition on the working model ${\cal M}_{w}$ would hold if the model $\bar{{\cal Q}}_{w}$ (or $\mathcal{Q}_{w}$, for $\Psi^{\#}$) falls with probability tending to one in a nice class such as the class of càdlàg functions with a universal bound on the sectional variation norm.